OpenStudy (anonymous):

(4 7th root of x)(6 cubed root of x)

OpenStudy (anonymous):

Simplify the expression and express the answer using a rational exponent. Assume that x denotes a positive number.

OpenStudy (anonymous):

\[(4\sqrt[7]{x})(6\sqrt[3]{x})\]\[=(4x^{\frac{1}{7}})(6x^{\frac{1}{3}})\]\[=24x^{\frac{1}{7}+\frac{1}{3}}\]\[=24x^{\frac{3}{21}+\frac{7}{21}}\]\[=24x^{\frac{10}{21}}\]